# Deep Learning for Cross-Domain Data Fusion in Urban Computing: Taxonomy, Advances, and Outlook

### v1

@@ 4.3 Alignment-based Data Fusion
Alignment fusion ensures that semantically related content across modalities is effectively combined. As illustrated in Figure 12, methods for multi-modal alignment span categories such as cross-modal attention mechanism (Sec.4.3.1) and multi-modal encoder-based fusion (Sec.4.3.2), enabling understanding and collaboration between disparate sources of information.
@@ 4.3.1 Attention-based Alignment
Attention mechanism Vaswani et al. (2017), especially multi-modal cross-attention is a fusion technique crucial for integrating information across diverse modalities, such as text and images Wei et al. (2020); Ji et al. (2020); Yang et al. (2023c). The cross-attention mechanism can be summarized in the following three steps. First, project the feature vectors of modalities X and Y into query (Q), key (K), and value (V) spaces using learnable parameter matrices $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$, respectively:
@@
In recent years, the urban computing community leveraged such fusion mode and its variants for comprehensive urban modality alignment. For example, to model the relations among the target road attributes, Yin et al. (2020) generated task-specific fused representations by applying attention-based feature fusion of location, bearing, speed, and map context. Zhang et al. (2021b) applied a GAT-based attention mechanism in learning region representations from two views of the built correlations (i.e., human mobility view and region attribute view), and a joint learning module to fuse multi-view embeddings. In this proposed multi-view joint learning module shown in Figure 13, the self-attention layer enables information sharing across all views and the fusion layer is responsible for combining multi-view representations via adaptive weights.
Figure 13: The architecture of multi-view joint learning module, consisting of a self-attention layer and a fusion layer Zhang et al. (2021b).
The MVMT-STN model Wang et al. (2023c) included a multi-view GCN to capture the global semantic dependency between POI, road network, and risk information. Huang et al. (2022a) used a transformer-based aggregation layer to model the graph structure containing both toponym and spatial knowledge. Qiang et al. (2023) designed a transformer encoder to represent the features of each package by integrating those of other candidate packages. The RankETPA model Wen et al. (2023b) was designed for package pick-up arrival time prediction, and there is an attention module to model the interaction between the package’s features and the courier’s features.
The other similar attentional fusion work includes Xi et al. (2022) (POI-view and geographic-view representations), Yuan et al. (2021a) (region-level and inter-traffic correlations), Wang et al. (2021a) (geographical and semantic spatio-temporal representations) and Yan et al. (2023) (satellite visual and textual representations).
@@ 4.3.2 Encoder-based Alignment
Encoder-based fusion entails the integration of multiple modalities into a shared encoder architecture, as opposed to employing separate encoders for each modality. This approach leverages a unified deep learning model (e.g., RNN variants Song et al. (2016); Jiang et al. (2019) and self-attention methods Cai et al. (2023); Li et al. (2023d)) to collectively process and extract meaningful representations from diverse sources of information. By cohesively injecting multi-modal data into a single encoder, the model is inherently encouraged to capture intricate inter-modal relationships and dependencies.
For example, Song et al. (2016) developed the DeepTransport model for mobility simulation and transportation mode prediction, where heterogeneous data including GPS records and transportation information are fed into a deep LSTM learning architecture. Such multi-layer LSTM has been demonstrated to be able to learn at different time scales over the input Hermans and Schrauwen (2013). The DeepUrbanEvent system Jiang et al. (2019) contained a ConvLSTM Shi et al. (2015) encoder module for simultaneous multi-step forecasting of crowd density and crowd flow. The ConvLSTM extends the fully connected LSTM (FC-LSTM) to have convolutional structures in both input-to-state and state-to-state transitions. The RegionDCL framework Li et al. (2023d) leveraged the Transformer encoder with average-pooling as the region-level encoder of building features and POI information. Cai et al. (2023) proposed a multi-level graph encoder equipped with a GAT encoding module to capture couriers’ both high-level transfer modes between AOIs and low-level transfer models between locations.
@@ 4.4 Contrast-based Data Fusion
Contrastive learning, a pivotal paradigm in machine learning, can be categorized based on contrast creation Le-Khac et al. (2020); Nakada et al. (2023). The categorization encompasses methods such as instance contrast, batch contrast, and temporal contrast, which each focuses on contrasting representations through augmentations, batch comparisons, and temporal shifts, respectively.
The InfoNCE (Noise-Contrastive Estimation) loss, a common objective function in contrastive learning, aims to maximize the similarity between positive pairs and minimize the similarity between negative pairs Oord et al. (2018). Mathematically, the InfoNCE loss is formulated as follows:
@@
CLIP (Contrastive Language-Image Pre-training) is a typical model in contrastive learning, designed to concurrently learn representations of images and text Radford et al. (2021). CLIP can be considered a form of instance contrast because it encourages the model to align representations of a given instance across different modalities. The model achieves this by minimizing the negative logarithmic probability of similarity between positive image-text pairs while contrasting against negative pairs. Mathematically, the CLIP loss function ($\mathcal{L}_{\text{CLIP}}$) is expressed as:
@@ 4.4.1 Instance Contrast-based Fusion
In instance contrast, the model aims to fuse the knowledge between different views of the same instance. It is effective for capturing intricate details within individual data points, promoting the model’s ability to recognize fine-grained patterns and features Le-Khac et al. (2020); Deldari et al. (2022); Liu et al. (2021).
Inspired by the computer vision domain Xu et al. (2023d); Chengchuang et al. (2021); Zoph et al. (2020), urban multi-modal research started to construct self-augmented data as contrastive pairs. For example, Li et al. (2022b) explored self-similarity across urban images to construct contrastive samples via data augmentation methods including rotation, gray-scale, and flipping. Similarly, Bai et al. (2023) learned the physical properties of the geographies via intra-modal contrast among very high resolution (VHR) image augmentations. In addition to imagery augmentation, Zhang et al. (2023c) designed three POI-level augmentations for intra-view contrastive fusion: random insertion, random deletion, and random replacement. Furthermore, Mao et al. (2022) introduced domain-specific augmentations for road-road contrast and trajectory-trajectory contrast separately, i.e., road segment with its contextual neighbors and trajectory with its detour replaced and dropped alternatives.
@@ 4.4.2 Batch Contrast-based Fusion
Batch contrast involves contrasting samples within the same batch. It introduces a form of global context, where the model learns to distinguish features not just within instances but also in relation to the entire batch Le-Khac et al. (2020); Deldari et al. (2022); Liu et al. (2021).
Geographical similarity-guided contrastive learning is pivotal in urban computing because urban images adhere to Tobler’s First Law of Geography Miller (2004). For instance, Li et al. (2022b) further enhanced the contrastive learning method by taking into account geographical similarity and minimizing the feature distance between two images that are geo-adjacent.
Besides, the CLIP-based paradigm has been explored in urban computing in recent years. For instance, Liu et al. (2023g) proposed the KnowCL model for socioeconomic prediction, which is the first solution that introduces the regional knowledge graph-based semantics and its associated imagery representation as contrastive pairs. Except for self-augmented contrast, Bai et al. (2023) also bridged the socio-economic semantic gap through inter-modal contrast between VHR images and POIs. The RecMVC model of Zhang et al. (2023c) included an inter-view contrastive learning module between POI and mobility, serving as a soft co-regularizer to transfer knowledge across multi-views. Liu et al. (2022a) designed a trajectory contrastive learning paradigm, where hub and link representations in the same trajectory are enforced to have a higher correlation with one another. Likewise, Mao et al. (2022) introduced road-trajectory cross-scale contrast to bridge the two scales by maximizing the total mutual information. This contrast is elaborately tailored via novel positive sampling and adaptive weighting strategies. Following the conventional CLIP setting, Yan et al. (2023) leveraged satellite imagery and associated LLM-generated description as positive pairs, to learn a robust visual embedding for urban region profiling (depicted in Figure 15).
Figure 15: The framework of UrbanCLIP, the first-ever LLM-enhanced framework that integrates the knowledge of textual modality into urban imagery profiling. It utilizes the CLIP rationale for language-image pertaining Yan et al. (2023).
@@ 4.5 Generation-based Data Fusion
The primary objective of a generation-based model is to produce the desired data based on specified input conditions. Both input and output data may manifest in diverse formats Yan et al. (2023); Yuanshao et al. (2023); Zhang et al. (2023a), rendering it a favorable choice for data fusion. Compared to the general Computer Vision (CV) or Natural Language Processing (NLP) area, the field of urban computing exhibits a relatively limited volume of generative research contributions. However, in recent years, there has been a growing number of generative works emerging, particularly influenced by the rise of LLMs Liu et al. (2023d). Based on the specific methods of generation, here we categorize the generation-based data fusion into four types, including autoregressive (Sec.4.5.1), mask modeling (Sec.4.5.2), diffusion-based (Sec.4.5.3), and LLM-enhanced models (Sec.4.5.4).
@@ 4.5.1 Autoregressive Model
Gradually, the definition of input data and output data for autoregressive model has expanded from text to encompass multi-modal data. Owing to the multi-modal nature of the attention mechanism Vaswani et al. (2017), whereby queries, keys, and values can stem from different modalities, it has facilitated the exploration of multi-modal fusion. VirTex Desai and Johnson (2021) proposed that compared with contrastive learning which uses classification labels as a learning signal, captions can provide a more semantically dense learning signal. SimVLM Wang et al. (2021f) reformulated the typical encoder-decoder architecture for end-to-end training with a single prefix language modeling objective, achieving the efficient utilization of weakly aligned image-text pairs. CoCa Yu et al. (2022) is a follow-up work to ALBEF Li et al. (2021a) and SimVLM Wang et al. (2021f), which creatively integrated attentional pooling methods in a creative manner, combining contrastive and generative approaches to achieve exceptional performance.
In the field of urban computing Zheng et al. (2014a), the generative decoder has played a significant role, even before the era of transformer models Vaswani et al. (2017). GeoMAN Liang et al. (2018) first introduced a multi-layer attention mechanism for spatio-temporal data prediction based on the encoder-decoder architecture. The generative decoder combined LSTM and Temporal Attention to predict the future performance of sensors. It achieved excellent performance in applications such as water quality prediction and air quality. Zhao et al. (2021b) implemented a bottom-up and top-down framework to do street-view image classification, which utilized RNN units to make predictions based on visual elements and contextual information. Besides, in urban computing area Zheng et al. (2014a), the limited availability and diverse nature of domain data, as well as the challenges in its acquisition, have led to increased interest in multi-modal pretraining in recent years.
@@ 4.5.2 Mask Modeling-based Fusion
The concept of masking modeling task has its origins in the field of NLP, contributing to the success of BERT Devlin et al. (2018). MAE He et al. (2022) is the first work that introduces masking modeling into CV. Compared with masking modeling works in NLP, MAE He et al. (2022) has a larger mask ratio due to the fact that the images have less information density than natural language He et al. (2022). GraphMAE Hou et al. (2022) unleashes the power of mask modeling for graph structure. Different from most graph autoencoders’ efforts in structure reconstruction, it proposes to focus on feature reconstruction with both a masking strategy and scaled cosine error. In the context of mask modeling structure for multi-modal data fusion, various modalities are typically provided as input, followed by the masking of a portion or all of the modal data at varying proportions. Subsequently, information from other modalities is integrated to reconstruct the masked data Kwon et al. (2022); Geng et al. (2022); Ding et al. (2023).
@@
MGeo Ding et al. (2023) combined text and geolocation to implement location embedding for query-POI matching, which utilized two unimodal masked language modeling (MLM) loss and one Multi-Modal MLM loss to enforce information interaction. As illustrated in Figure 16, G2PTL Wu et al. (2023a) is a graph-based pre-trained model for text-based delivery address embedding. It leveraged MLM as one of the pre-training tasks to simulate input noise in missing or incorrect address cases in real-world scenarios. QUERT Xie et al. (2023) adapted pre-trained language models to specific Travel Domain Search applications, which designed a Geography-aware masking strategy to force the pre-trained model to pay more attention to the geographical location phrases. ERNIE-GeoL Huang et al. (2022a) was dedicated to integrating toponym knowledge and spatial knowledge. By utilizing a Masked Language Modeling training procedure, it can acquire four types of toponym knowledge, including relationships between POI name, address, and type.
@@
Figure 16: The framework of G2PTL, a Geography-Graph Pre-Trained model for delivery address in the Logistics field, which uses the MLM task to learn the semantic information in the address Wu et al. (2023a).
@@
In the remote sensing area, despite the similarities between remote sensing images and RGB images, they also possess distinctions, with remote sensing images containing geographical information and temporal tags, as well as spectral multiple channels, and diverse resolutions and scales. SatMAE Cong et al. (2022) discussed the domain-specific temporal and spectral characteristics of satellite imagery, and device’s temporal encoding and spectral positional encoding respectively to adapt MAE structure to satellite imagery representation learning. Scale-MAE Reed et al. (2023) was designed to address another challenge in remote sensing satellite imagery: the variation in image scale due to data from multi-scale sensors. The pretraining process aims to learn a better representation of multiscale tasks by reconstructing low and high frequency features at different scales.

### v2

@@ 4.3 Alignment-based Data Fusion
Alignment fusion ensures that semantically related content across modalities is effectively combined. As illustrated in Figure 13, methods for multi-modal alignment span categories such as cross-modal attention mechanism (Sec.4.3.1) and multi-modal encoder-based fusion (Sec.4.3.2), enabling understanding and collaboration between disparate sources of information. Alignment-based approaches can achieve more precise modal alignment and exhibit flexibility, making them suitable for various general scenarios. However, they generally require higher computational resources, which makes them more appropriate for offline tasks in transportation and urban planning.
@@ 4.3.1 Attention-based Alignment
Attention mechanism [268], especially multi-modal cross-attention is a fusion technique crucial for integrating information across diverse modalities, such as text and images [296, 122, 327]. The cross-attention mechanism can be summarized in the following three steps. First, project the feature vectors of modalities X and Y into query (Q), key (K), and value (V) spaces using learnable parameter matrices $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$, respectively:
@@
In recent years, the urban computing community leveraged such fusion mode and its variants for comprehensive urban modality alignment. For example, to model the relations among the target road attributes, Yin et al. [339] generated task-specific fused representations by applying attention-based feature fusion of location, bearing, speed, and map context. Zhang et al. [368] applied a GAT-based attention mechanism in learning region representations from two views of the built correlations (i.e., human mobility view and region attribute view), and a joint learning module to fuse multi-view embeddings. In this proposed multi-view joint learning module shown in Figure 14, the self-attention layer enables information sharing across all views and the fusion layer is responsible for combining multi-view representations via adaptive weights.
Figure 14: The architecture of multi-view joint learning module, consisting of a self-attention layer and a fusion layer[368]. $\mathcal{E}_{i}^{\prime}$ is the representation for the $i$-th view of global information, and $\alpha$ is the weight of global information.
The MVMT-STN model [288] included a multi-view GCN to capture the global semantic dependency between POI, road network, and risk information. Huang et al. [110] used a transformer-based aggregation layer to model the graph structure containing both toponym and spatial knowledge. Qiang et al. [225] designed a transformer encoder to represent the features of each package by integrating those of other candidate packages. The RankETPA model [298] was designed for package pick-up arrival time prediction, and there is an attention module to model the interaction between the package’s features and the courier’s features.
The other similar attentional fusion work includes [309] (POI-view and geographic-view representations), [349] (region-level and inter-traffic correlations), [272] (geographical and semantic spatio-temporal representations) and [323] (satellite visual and textual representations).
@@ 4.3.2 Encoder-based Alignment
Encoder-based fusion entails the integration of multiple modalities into a shared encoder architecture, as opposed to employing separate encoders for each modality. This approach leverages a unified deep learning model (e.g., RNN variants [248, 123] and self-attention methods [28, 156]) to collectively process and extract meaningful representations from diverse sources of information. By cohesively injecting multi-modal data into a single encoder, the model is inherently encouraged to capture intricate inter-modal relationships and dependencies.
For example, Song et al. [248] developed the DeepTransport model for mobility simulation and transportation mode prediction, where heterogeneous data including GPS records and transportation information are fed into a deep LSTM learning architecture. Such multi-layer LSTM has been demonstrated to be able to learn at different time scales over the input [103]. The DeepUrbanEvent system [123] contained a ConvLSTM [243] encoder module for simultaneous multi-step forecasting of crowd density and crowd flow. The ConvLSTM extends the fully connected LSTM (FC-LSTM) to have convolutional structures in both input-to-state and state-to-state transitions. The RegionDCL framework [156] leveraged the Transformer encoder with average-pooling as the region-level encoder of building features and POI information. Cai et al. [28] proposed a multi-level graph encoder equipped with a GAT encoding module to capture couriers’ both high-level transfer modes between AOIs and low-level transfer models between locations.
@@ 4.4 Contrast-based Data Fusion
Contrastive learning, a pivotal paradigm in machine learning, can be categorized based on contrast creation [147, 202]. The categorization encompasses methods such as instance contrast, batch contrast, and temporal contrast, which each focuses on contrasting representations through augmentations, batch comparisons, and temporal shifts, respectively. Compared to alignment-based methods, the contrast-based data fusion method enhances discrimination through negative sample augmentation. However, it imposes stringent requirements on the selection of negative samples and batch size. Due to its demand for large datasets, it is well-suited for downstream tasks where data is abundant or easily accessible, such as those related to transportation, environment and economy.
The InfoNCE (Noise-Contrastive Estimation) loss, a common objective function in contrastive learning, aims to maximize the similarity between positive pairs and minimize the similarity between negative pairs [210]. Mathematically, the InfoNCE loss is formulated as follows:
@@
CLIP (Contrastive Language-Image Pre-training) is a typical model in contrastive learning, designed to concurrently learn representations of images and text [227]. CLIP can be considered a form of instance contrast because it encourages the model to align representations of a given instance across different modalities. The model achieves this by minimizing the negative logarithmic probability of similarity between positive image-text pairs while contrasting against negative pairs. Mathematically, the CLIP loss function ($\mathcal{L}_{\text{CLIP}}$) is expressed as:
@@ 4.4.1 Instance Contrast-based Fusion
In instance contrast, the model aims to fuse the knowledge between different views of the same instance. It is effective for capturing intricate details within individual data points, promoting the model’s ability to recognize fine-grained patterns and features [147, 56, 183].
Inspired by the computer vision domain [319, 46, 407], urban multi-modal research started to construct self-augmented data as contrastive pairs. For example, Li et al. [154] explored self-similarity across urban images to construct contrastive samples via data augmentation methods including rotation, gray-scale, and flipping. Similarly, Bai et al. [10] learned the physical properties of the geographies via intra-modal contrast among very high resolution (VHR) image augmentations. In addition to imagery augmentation, Zhang et al. [366] designed three POI-level augmentations for intra-view contrastive fusion: random insertion, random deletion, and random replacement. Furthermore, Mao et al. [196] introduced domain-specific augmentations for road-road contrast and trajectory-trajectory contrast separately, i.e., road segment with its contextual neighbors and trajectory with its detour replaced and dropped alternatives.
@@ 4.4.2 Batch Contrast-based Fusion
Batch contrast involves contrasting samples within the same batch. It introduces a form of global context, where the model learns to distinguish features not just within instances but also in relation to the entire batch [147, 56, 183].
Geographical similarity-guided contrastive learning is pivotal in urban computing because urban images adhere to Tobler’s First Law of Geography [198]. For instance, Li et al. [154] further enhanced the contrastive learning method by taking into account geographical similarity and minimizing the feature distance between two images that are geo-adjacent.
Besides, the CLIP-based paradigm has been explored in urban computing in recent years. For instance, Liu et al. [186] proposed the KnowCL model for socioeconomic prediction, which is the first solution that introduces the regional knowledge graph-based semantics and its associated imagery representation as contrastive pairs. Except for self-augmented contrast, Bai et al. [10] also bridged the socio-economic semantic gap through inter-modal contrast between VHR images and POIs. The RecMVC model of Zhang et al. [366] included an inter-view contrastive learning module between POI and mobility, serving as a soft co-regularizer to transfer knowledge across multi-views. Liu et al. [173] designed a trajectory contrastive learning paradigm, where hub and link representations in the same trajectory are enforced to have a higher correlation with one another. Likewise, Mao et al. [196] introduced road-trajectory cross-scale contrast to bridge the two scales by maximizing the total mutual information. This contrast is elaborately tailored via novel positive sampling and adaptive weighting strategies. Following the conventional CLIP setting, Yan et al. [323] leveraged satellite imagery and associated LLM-generated description as positive pairs, to learn a robust visual embedding for urban region profiling (depicted in Figure 16). Subsequent work, UrbanVLP [97], has broadened the contrastive learning paradigm to encompass multi-granularity visual clues, including satellite and street-view images. Concurrently, it also proposes effective methods to guarantee the generated text quality.
Figure 16: The framework of UrbanCLIP, the first-ever LLM-enhanced framework that integrates the knowledge of textual modality into urban imagery profiling. It utilizes the CLIP rationale for language-image pertaining [323].
@@ 4.5 Generation-based Data Fusion
The primary objective of a generation-based model is to produce the desired data based on specified input conditions. Both input and output data may manifest in diverse formats [323, 356, 358], rendering it a favorable choice for data fusion. By generating new content that correlates with input data, generative models are driven to discern intricate correspondences between multimodal information, thereby facilitating efficient information aggregation. However, these models are marked by significant computational complexity and considerable training challenges. Consequently, they are not suitable for tasks requiring high real-time performance. Instead, they excel with large datasets [131], making them particularly well-suited for industries such as transportation, economics, and energy. Compared to the general Computer Vision (CV) or Natural Language Processing (NLP) area, the field of urban computing exhibits a relatively limited volume of generative research contributions. However, in recent years, there has been a growing number of generative works emerging, particularly influenced by the rise of LLMs [180]. Based on the specific methods of generation, here we categorize the generation-based data fusion into four types, including autoregressive (Sec.4.5.1), mask modeling (Sec.4.5.2), diffusion-based (Sec.4.5.3), and LLM-enhanced models (Sec.4.5.4).
@@ 4.5.1 Autoregressive Model
Gradually, the definition of input data and output data for autoregressive model has expanded from text to encompass multi-modal data. Owing to the multi-modal nature of the attention mechanism [268], whereby queries, keys, and values can stem from different modalities, it has facilitated the exploration of multi-modal fusion. VirTex [57] proposed that compared with contrastive learning which uses classification labels as a learning signal, captions can provide a more semantically dense learning signal. SimVLM [295] reformulated the typical encoder-decoder architecture for end-to-end training with a single prefix language modeling objective, achieving the efficient utilization of weakly aligned image-text pairs. CoCa [344] is a follow-up work to ALBEF [152] and SimVLM [295], which creatively integrated attentional pooling methods in a creative manner, combining contrastive and generative approaches to achieve exceptional performance.
In the field of urban computing [391], the generative decoder has played a significant role, even before the era of transformer models [268]. GeoMAN [162] first introduced a multi-layer attention mechanism for spatio-temporal data prediction based on the encoder-decoder architecture. The generative decoder combined LSTM and Temporal Attention to predict the future performance of sensors. It achieved excellent performance in applications such as water quality prediction and air quality. Zhao et al. [382] implemented a bottom-up and top-down framework to do street-view image classification, which utilized RNN units to make predictions based on visual elements and contextual information. Besides, in urban computing area [391], the limited availability and diverse nature of domain data, as well as the challenges in its acquisition, have led to increased interest in multi-modal pretraining in recent years.
@@ 4.5.2 Mask Modeling-based Fusion
The concept of masking modeling task has its origins in the field of NLP, contributing to the success of BERT [60]. MAE [102] is the first work that introduces masking modeling into CV. Compared with masking modeling works in NLP, MAE [102] has a larger mask ratio due to the fact that the images have less information density than natural language [102]. GraphMAE [104] unleashes the power of mask modeling for graph structure. Different from most graph autoencoders’ efforts in structure reconstruction, it proposes to focus on feature reconstruction with both a masking strategy and scaled cosine error. In the context of mask modeling structure for multi-modal data fusion, various modalities are typically provided as input, followed by the masking of a portion or all of the modal data at varying proportions. Subsequently, information from other modalities is integrated to reconstruct the masked data [142, 83, 61].
@@
MGeo [61] combined text and geolocation to implement location embedding for query-POI matching, which utilized two unimodal masked language modeling (MLM) loss and one Multi-Modal MLM loss to enforce information interaction. As illustrated in Figure 17, G2PTL [304] is a graph-based pre-trained model for text-based delivery address embedding. It leveraged MLM as one of the pre-training tasks to simulate input noise in missing or incorrect address cases in real-world scenarios. QUERT [314] adapted pre-trained language models to specific Travel Domain Search applications, which designed a Geography-aware masking strategy to force the pre-trained model to pay more attention to the geographical location phrases. ERNIE-GeoL [110] was dedicated to integrating toponym knowledge and spatial knowledge. By utilizing a Masked Language Modeling training procedure, it can acquire four types of toponym knowledge, including relationships between POI name, address, and type.
@@
Figure 17: The framework of G2PTL, a Geography-Graph Pre-Trained model for delivery address in the Logistics field, which uses the MLM task to learn the semantic information in the address [304].
@@
In the remote sensing area, despite the similarities between remote sensing images and RGB images, they also possess distinctions, with remote sensing images containing geographical information and temporal tags, as well as spectral multiple channels, and diverse resolutions and scales. SatMAE [49] discussed the domain-specific temporal and spectral characteristics of satellite imagery, and device’s temporal encoding and spectral positional encoding respectively to adapt MAE structure to satellite imagery representation learning. Scale-MAE [230] was designed to address another challenge in remote sensing satellite imagery: the variation in image scale due to data from multi-scale sensors. The pretraining process aims to learn a better representation of multiscale tasks by reconstructing low and high frequency features at different scales.